OpenStudy (anonymous):

simplify the following expressions 1. 2a + 3b - a 2. (4x + 3y) - (2x -5y) 3. -3(2x - 1) – 5 4.7xy - 2xz + 4xz - 12xyz + 2zx 5.-3(2x - 5) - 4(x + 7)

OpenStudy (anonymous):

can someone please help me

OpenStudy (anonymous):

1. a+3b

OpenStudy (anonymous):

2. 2x+8y

OpenStudy (skullpatrol):

Do you recall the distributive property?

OpenStudy (skullpatrol):

$$a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

@Behindthemask are you still there?

OpenStudy (skullpatrol):

If you don't understand something you need to ask questions.

OpenStudy (anonymous):

yes.

OpenStudy (skullpatrol):

$$a(b-c)=ab-ac $$ $$-3(2x - 1) – 5$$ $$-3(2x)-(-3)(1)-5$$ $$-6x-(-3)-5$$ $$-6x+3-5$$ $$-6x-2$$

OpenStudy (anonymous):

@skullpatrol can you help with the other two remaining as well

OpenStudy (anonymous):

ill do one and you have to do the other

OpenStudy (anonymous):

for number 4

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

like terms are -2xz+4xz+2xz=4xz

OpenStudy (anonymous):

finish it